OpenStudy (anonymous):

(_ _)^2 = (sec x-1)(sec x+1)

OpenStudy (anonymous):

(sec x-1)(sec x+1) = sec^2x-1 tan^2x = sec^2x-1 identity ,.. http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

(tanx)^2

OpenStudy (anonymous):

thank you both!

OpenStudy (anonymous):

np man yw